OpenStudy (anonymous):

1. Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66.

OpenStudy (anonymous):

the nth term in a AP are given by: a+(n-1)d where a is the first term and d is the common difference. Here we have the first term being -12 ie a = -12. 27th term is 66, so we have: -12+(27-1)d=66 26d=66+12=78 d=78/26=3 Therefore the nth term is: -12+3(n-1) So the 42nd term is: -12+3(42-1)=111

OpenStudy (goformit100):

A Warm Welcome To Open Study